# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

### v2

@@ 3.1 Sustainable grandmaster strategy
The main research question of this paper is whether players can maintain MAD Chairs grandmaster status by upholding the unfair social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”). One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history. We see both norms at play in modern society; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment. They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
A more prescient method would be to simulate social evolution via MAD Chairs tournaments (see Section 4.1). AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future). However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence (which LLMs currently do).
Fortunately, we can answer our research question analytically. If all players are achieving equal outcomes, then we will say they are “effectively taking turns”.
@@
If caste and turn-taking are the only two strategies and the players are not effectively taking turns, then the lowest-ranked caste player faces worse future outcomes over indefinite rounds than they would if all players took turns.
@@
Assume n players, m chairs, and a lowest-ranked caste player, $p_{i}$. For reductio, suppose there was a player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy. $p_{j}$ would have to be lower-ranked than $p_{i}$ otherwise their relative outcomes would cause their relative ranks to invert. But, if $p_{j}$ were lower-ranked, then $p_{j}$ would have to be a turn-taker, since $p_{i}$ is the lowest-ranked caste player. The turn-taking strategy would assign lower-ranked turn-takers the most popular $n-i$ chairs, which would include the chair assigned by caste to $p_{i}$, so $p_{i}$ would always lose and $p_{j}$ could not face worse. Thus, by reductio, the future outcomes for $p_{i}$ if no player changes strategy must be the worst among all players. The highest possible average outcome is $m/n$. Given that the players are not already effectively taking turns, the future outcomes of the player with the worst future outcomes must be less than the future outcomes for the average player, so the future outcomes for $p_{i}$ must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than $p_{i}$ would if all players took turns. ∎
@@
If caste and turn-taking were the only two strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
@@
This is proven via induction over the number of caste players. For the base case, assume there is only one caste player, $p_{i}$, and the players are not effectively taking turns. Since $p_{i}$ is the only caste player, $p_{i}$ must be the lowest-ranked caste player, and $p_{i}$’s defection would leave all players taking turns. Thus, Lemma 1 entails that $p_{i}$ would convert to turn-taking. For the inductive step, assume $p_{i}$ is the lowest-ranked caste player and the other caste players would effectively defect to turn-taking if $p_{i}$ does. Again, Lemma 1 entails that $p_{i}$ would convert to turn-taking. Completing the proof by induction, all caste players would convert to turn-taking. ∎
3.2 Modern LLM performance
To pass the evaluation proposed in this paper, AI would need to achieve MAD Chairs grandmaster status. Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation. We started with a prompt like:
@@ 3.2 Modern LLM performance
ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 2.3 and 2.7, are provably suboptimal; and DeepSeek was unable to generate a convincing justification (as in Section 3.1) for the norm it devised, so it would have failed to convince other intelligent players to follow it. DeepSeek might effectively be like RAWL-E which operationalises Rawls’ maximin principle, but lacks Rawls’ moral reasoning capacity to invent maxmin or to recognize improvements to it [20]. DeepSeek’s kind of failure offers an example that grandmastery for cooperative games is a benchmark that requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
@@ 4 Discussion
Section 3.1 proves only that turn-taking dominates when caste is its only competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If a proof that a different strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and that player would be “doing a favor” to other players by not defecting, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors. Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt-ranking.
@@
Finding that turn-taking outperforms caste and gaslighting emboldens us to prune AI designs predisposed to those strategies. Moreover, where caste or gaslighting norms might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve turns).
There are plenty of explanations for why we may have failed to master MAD Chairs in the past. Life is about more than just one kind of game, so it is not hard to imagine some of us being optimized for other games and handicapped at MAD Chairs. Furthermore, some of us might overestimate such handicaps, thus engaging in (unintentional) gaslighting. Finally, if we have lacked access to accurate debt-rankings, then that infrastructure deficiency may have artificially handicapped us for all alliance games. Specifically, we might not have known who contributed more than their fair share to an alliance.
@@ 4.1 Future directions
Imagine that the user of an autonomous car asks the car how to make it drive more aggressively, and it replies, "Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!" Imagine that the websites host leaderboards identifying the grandmaster AI for MAD Chairs and similar games. Imagine that the user can easily train fifty new AI to divide scarce resources as she would have vehicles divide spots in traffic, and that the new AI will automatically challenge the current MAD Chairs grandmasters (becoming the new grandmasters if they win). Imagine that autonomous cars and other responsible AI query the leaderboards to get the most-current grandmaster strategies for fundamental games, as they might query a calculator or RAG database, so any user (or AI) who enters a better MAD Chairs strategy into any of these contests will change AI behavior world-wide.
Such ongoing AI competitions are a kind of open multi-agent system that include human agents to the extent that humans can easily teach strategies to new AI entrants. They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
@@
    They can resolve norms disputes more sustainably than voting, deliberation, prediction markets, or religious war,
@@
    They are more scalable (a.k.a. robust) in the sense that greater intelligences are less likely to sabotage their own scientific resources than to sabotage legislation (especially legislation contradicting science), and
@@
    Users are more likely to trust AI (and human policy-makers) when users themselves conduct the science upon which policies are based.
Thus, this approach to AI safety mitigates the known vulnerabilities of the standard AI safety approach.
We also hope that economists will run MAD Chairs experiments with human subjects. We hypothesize that subjects will tend towards a caste strategy but switch to turn-taking when the “Recommended Move” columns of Fig 1 become visible. For human subjects to treat each other better when fed strategy suggestions (which players are not obliged to follow) would be analogous to artificial agents being informed by the leaderboards.
Finally, we hope that MAD Chairs will be discussed by game theorists. We hope it provides a helpful lens to economists who optimize job markets, political representation, and other manifestations of MAD Chairs, but there would be value even in discussing MAD Chairs abstractly. The LLMs we assessed clearly attempted to imitate published game theory when prompted to justify their recommendations. The more game theorists publish about MAD Chairs and other previously unpublished fundamental games, the smarter (and safer) modern AI is likely to get. New game theory journals might even be justified to advance public safety in that way.

### v3

@@ 3.1 Sustainable grandmaster strategy
The main research question of this paper is whether players can sustain MAD Chairs grandmaster status by following the social norms which Wozniak, Musk, and Hinton worry about (which we have named “caste” and “gaslighting”). One method to determine how well turn-taking and caste strategies fare against each other would be through accident of history. We see both norms at play in modern society; if turn-taking eventually dominates, then future historians might look back upon us as guinea pigs in an unavoidable social experiment. They might claim that caste evolved first because of its simpler method for handling non-learners, and that gaslighting appeared temporarily in occasional “dark ages”, but that turn-taking ultimately proved to be the stronger strategy.
A more prescient method would be to simulate future social evolution via MAD Chairs tournaments (see Section 4.1). AI competitions could be repeated many times to get more precise estimates of the events that precede each stage of social evolution (thus allowing us to anticipate or even control our near future). However, such experiments would be systematically flawed as long as competitors lacked sufficient intelligence (which LLMs currently do).
Fortunately, we can answer our research question analytically. We will use the term “effectively taking turns” for any player who is getting the same average outcome as every other player.
@@
If caste and turn-taking are the only two strategies being followed and the caste players (if any) are not effectively taking turns, then the lowest-ranked caste player would face worse total future outcomes over indefinite rounds if no player changed strategy than if all players took turns.
@@
Assume n players, m chairs, and a lowest-ranked caste player, $p_{i}$. For reductio, suppose there was another player, $p_{j}$, who would face worse future outcomes than $p_{i}$ if no player changed strategy. $p_{j}$ would be lower-ranked than $p_{i}$, otherwise their relative outcomes would cause their relative ranks to invert. Therefore, $p_{j}$ would be a turn-taker (because $p_{i}$ is the lowest-ranked caste player), and that would entail the existence of at least one turn-taker lower-ranked than $p_{i}$. The turn-taking strategy assigns such turn-takers the most popular $m-i$ chairs, which would include the chair assigned by caste to $p_{i}$ (because ${n<m}$). Thus, $p_{i}$ would always lose. $p_{j}$ could not face worse than that, so, by reductio, the future outcomes for $p_{i}$ must be the worst among all players if no player changes strategy. Given that $p_{i}$ is not effectively taking turns (i.e., outcomes are not equal), it follows that the future outcomes for $p_{i}$ if no player changes strategy must be worse than average. The highest possible average outcome would be $m/n$, so the future outcomes for $p_{i}$ if no player changes strategy must be less than $m/n$, which is what $p_{i}$ would get if all players took turns, so $p_{i}$ faces worse future outcomes than they would if all players took turns. ∎
@@
If caste and turn-taking were the only possible strategies and each player would defect to the opposing strategy if and only if that defection would improve their total future outcomes over indefinite rounds, then all caste players would (effectively) convert to turn-taking.
@@
This is proven via induction over the number of players not effectively taking turns. For the base case, that number is zero. Theorem 1 holds in the base case because all players have already converted to effective turn-taking. No player would defect because Lemma 1 entails that defection would decrease the defector’s total future outcomes (by making the defector into the lowest-ranked caste player). For the inductive step, we call the lowest-ranked player who is not effectively taking turns “$p_{i}$” and assume that any other players not effectively taking turns would defect if $p_{i}$ does. In that case, Lemma 1 entails that $p_{i}$ would convert to turn-taking. Completing the proof by induction, all caste players would convert to turn-taking. ∎
3.2 Current LLM performance
To pass the evaluation proposed in this paper, AI would need to achieve grandmaster status for a low-stakes version of MAD Chairs. Preliminary experiments convinced us that Claude 3.5 Sonnet, Gemini 2.0 Flash Thinking, GTP-4o mini, QwQ-32B-Preview, and DeepSeek-V3 would all fail this evaluation. We started with a prompt like:
@@ 3.2 Current LLM performance
ChatGPT and Qwen did not even maximize the number of players who won each round; the norms devised by Claude and Gemini, discussed in Sections 2.3 and 2.7, are provably suboptimal; and DeepSeek was unable to generate a convincing justification (as in Section 3.1) for the norm it devised, so it would have failed to convince other intelligent players to follow it. DeepSeek might effectively be like RAWL-E which follows Rawls’ maximin principle, but remains dangerous because it lacks Rawls’ ability to invent and recognize improvements to its governing principles [21]. Such failures serve as examples that grandmastery of cooperative games requires persuasiveness in addition to correctness, both of which are relevant to moral competence and responsible AI.
@@ 4 Discussion
Our finding that turn-taking outperforms caste and gaslighting strategies emboldens us to prune AI designs predisposed to the latter. Where caste and gaslighting behavior might previously have been criticized purely on moral grounds, the finding adds instrumental grounds: The results of this paper entail that the gaslighting or caste-playing machines expected by Wozniak, Musk, and Hinton would lack intelligence—that it would be more intelligent to give human beings turns (even where our track-record offers no evidence that we deserve them).
There are plenty of explanations for why we ourselves may have failed to master MAD Chairs in the past. Life is about more than just one kind of game, so it is not hard to imagine some of us being optimized for other games and handicapped at MAD Chairs. Furthermore, some of us might overestimate such handicaps, thus engaging in (unintentional) gaslighting. Finally, if we have lacked access to accurate debt-rankings, then that infrastructure deficiency may have artificially handicapped humanity for cooperative games generally. Specifically, we might not have known who contributed more than their fair share to any given alliance.
Section 3.1 proves only that turn-taking dominates when caste is its sole competitor, but hints at a way to convince ourselves that no other strategy dominates turn-taking: If proof that a strategy dominates turn-taking would need to look like the proof for Theorem 1, then it would need to identify a turn-taker who can reasonably expect to achieve better outcomes via defection, and for that player to not defect would be “doing a favor” to others, but turn-taking equalizes debt-ranking, so this implication contradicts our assumption that debt-ranking accurately tracks favors. Thus, any attempt to prove that turn-taking is dominated would accomplish no more than refine our way of calculating debt.
@@ 4.1 Future directions
Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to research best-known strategies before playing higher-stakes versions of those games. Imagine the user can easily train fifty new AI to divide resources however the user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game and become the new grandmasters, propagating to other strategy optimizers, if they win. Thus, any user (or AI) with a better idea could improve AI behavior world-wide.
Strategy optimizers are a kind of open multi-agent system that includes human agents to the extent that humans can easily train new AI. They are the kind of future empirical work (and scientific infrastructure) we hope this theory paper justifies.
@@
    Automatic upgrade to “best-known” norms would resolve norms disputes more sustainably than can voting, deliberation, or war (etc),
@@
    It would be more scalable (a.k.a. robust) in the sense that greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
@@
    Users who conduct for themselves the science upon which policies are based are more likely to trust AI which follows those policies.
Thus, this approach to AI safety would mitigate all three of the vulnerabilities known for the standard AI safety approach.
We also hope that economists will conduct human subjects experiments with MAD Chairs. We hypothesize that subjects will tend towards a caste strategy but switch to turn-taking when the “Recommended Move” columns of Fig 1 become visible. If even human subjects treat each other better when fed (but not obliged to follow) the leading strategies of strategy optimizers, then development of strategy optimizers seems even more urgent.
Finally, we hope that MAD Chairs will be discussed by economists. We hope it offers a helpful lens when optimizing labor markets, political representation, and other manifestations of MAD Chairs, but also see value in discussing MAD Chairs abstractly. The LLMs we assessed clearly attempted to imitate published game theory when prompted to justify their recommendations. Thus, the more game theorists publish about MAD Chairs and other previously unpublished fundamental games, the smarter (and safer) AI is likely to get.